### v4

@@ 7.2 MGS for Quasigroups
The proof of Lemma 7.9 shows that for any quasigroup $Q$ and any generating set $S\subseteq Q$, every element $g\in Q$ can be realized with a parse tree of depth $O(\log^{2}|Q|)$. In contrast, Wolf [Wol94] establishes the existence of parse trees of depth $O(\log|Q|)$.
For proving Theorem 7.7 we follow essentially the ideas of [Fle22] (though we avoid introducing the notion of Cayley circuits); however, we have to use a more refined approach by not simply guessing an SLP but the complete information given in Lemma 7.9. Fleischer obtained a $\mathsf{quasiAC^{0}}$ bound for Membership for groups by then showing that the Cayley circuits for this problem can be simulated by a depth-$2$ $\mathsf{quasiAC^{0}}$ circuit. We will instead directly analyze the straight-line programs using an $\exists^{\log^{2}n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ (sequential) algorithm.
Proof of Theorem 7.7.
@@
In the $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ computation we compute for all $i$ the elements $g_{i}^{(\mu)}=P(z_{0}z_{1}^{e_{i,1}^{(\mu)}}\cdots z_{i-1}^{e_{i,i-1}^{(\mu)}})$ and verify that $z_{i}$ is either $x_{\ell_{i}}$ or can be written as $ab$, $a\backslash b$ or $a/b$ where $a=g_{i}^{(1)}\backslash g_{i}^{(2)}$ and $b=g_{i}^{(3)}\backslash g_{i}^{(4)}$ – thus, verifying the conditions of Lemma 7.9. Finally, with the same technique we verify that $g\in L(k)$ by checking whether $g=P(z_{0}z_{1}^{f_{1}^{(1)}}\cdots z_{k}^{f_{k}^{(1)}})\backslash P(z_{0}z_{1}^{f_{1}^{(2)}}\cdots z_{k}^{f_{k}^{(2)}})$. ∎
Proof of Theorem 7.5.
(a) Let $G$ denote the input quasigroup (of order $n$). First, observe that every quasigroup has a generating set of size $\leq\lceil\log n\rceil$ [Mil78]. Therefore, we start by guessing a subset $X\subseteq G$ of size at most $\leq\lceil\log n\rceil$ (resp. the size bound given in the input). For this, we use $O(\log^{2}n)$ existentially quantified non-deterministic bits ($\exists^{\log^{2}n}$). Furthermore, we also guess some additional information, namely, the sequence $z_{1},\dots,z_{k}\in G$ from Lemma 7.9 and, just as in the proof of Theorem 7.7, the witnesses that this is, indeed, a cube-like sequence. As in the proof of Theorem 7.7 this takes again $O(\log^{2}n)$ existentially quantified bits.
In the next step, we verify whether $X$ actually generates $G$. This is done by checking for all $g\in G$ (universally verifying $O(\log n)$ bits, $\forall^{\log n}$) whether $g\in\langle X\rangle$, which can be done in $\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{quasiAC^{0}}$ just as in Theorem 7.7. Here it is crucial to note that, as we have already guessed the cube-like generating set in the outer existential block, it suffices to guess $O(\log n)$ bits for the exponents. Finally, as in the proof of Theorem 7.7, we verify in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ using the additional information guessed above that $z_{1},\dots,z_{k}\in G$ is indeed, a cube-like sequence. This concludes the proof of the bound (a) for the decision variant.
(b) To find a minimum-sized generating sequence, we can enumerate all possible generating sets $X$ of size at most $\log n$ in $\mathsf{quasiAC^{0}}\cap\mathsf{DSPACE}(\log^{2}n)$. If we want to compute the minimum generating set, we first, have to find a generating set $X$ and then we have to check whether $X$ is actually of smallest possible size. To do this, in a last step, we use the decision variant from above to check that there is no generating set of size at most $\left|\mathinner{X}\right|-1$ for $G$. ∎
@@
While it is possible to directly reduce MGS to Membership in $\mathsf{quasiAC^{0}}$, we obtain slightly better bounds by instead directly leveraging the Reachability Lemma 7.9. It is also possible to $\mathsf{quasiAC^{0}}$-reduce Quasigroup Isomorphism to Membership. This formalizes the intuition that, from the perspective of $\mathsf{quasiAC^{0}}$, membership testing is an essential subroutine for isomorphism testing and MGS. This might seem surprising, as in the setting of groups, Membership belongs to $\mathsf{L}$, while MGS belongs to $\textsf{AC}^{1}(\textsf{L})$ (Thm. 7.1), yet it is a longstanding open problem whether Group Isomorphism is in P.
@@ 7.3 MGS for Magmas
The decision variant of Minimum Generating Set is $\mathsf{NP}$-complete under many-one $\mathsf{AC^{0}}$ reductions for commutative magmas (unital or not).
This $\mathsf{NP}$-completeness result helps explain the use of Integer Linear Programming in practical heuristic algorithms for the search version of this problem in magmas, e. g., [JMV23].
For the closely related problem they called Log Generators—given the multiplication table of a binary function (=magma) of order $n$, decide whether it has a generating set of size $\leq\lceil\log_{2}n\rceil$—Papadimitriou and Yannakakis proved that Log Generators of Magmas is $\exists^{\log^{2}n}\mathsf{P}$-complete under polynomial-time reductions [PY96, Thm. 7]. Our proof uses a generating set of size roughly $\sqrt{n}$, where $n$ is the order of the magma; this is analogous to the situation that Log-Clique is in $\exists^{\log^{2}n}\mathsf{P}$, while finding cliques of size $\Theta(\sqrt{n})$ in an $n$-vertex graph is $\mathsf{NP}$-complete.
Proof.
@@
Aside from multiplication being commutative (e. g., we also have $X_{i}T=\overline{X}_{i}$, etc.), all other products are defined as 00.
The idea is that the presence of $S_{j,k}$ in a word indicates that clauses $j,j+1,\dotsc,k$ have been satisfied. This interpretation aligns with the multiplication above: viz. $C_{j}X=S_{j,j}$ if $X$ satisfies $C_{j}$. If clauses $j$ through $k$ are satisfied ($S_{j,k}$) and clauses $k+1$ through $\ell$ are satisfied ($S_{k+1,\ell}$), then in fact clauses $j$ through $\ell$ are satisfied ($S_{j,k}S_{k+1,\ell}=S_{j,\ell}$). We use “$S$” for all of these as a mnemonic for “satisfied” and also because $S=S_{1,m}$ acts as a “swap” on literals.
@@
Identity elements. We may add a new element $e$, relations $ea=ae=a$ for all $a$, and include $e$ as a generator. When this is done, the reduction queries whether the algebra is generated by $n+m+1$ elements or more than that many, since the element $e$ must be contained in every generating set. ∎
As with most $\mathsf{NP}$-complete decision versions of optimization problems, we expect that the exact version—given a magma $M$ and an integer $k$, decide whether the minimum generating set has size exactly $k$—is $\mathsf{DP}$-complete, but we leave that as a (minor) open question.
@@ 8 Conclusion
The analogous result is known for groups, by reducing to the connectivity problem on Cayley graphs. The best known bound for quasigroups is $\mathsf{SAC}^{1}$ due to Wagner [Wag10]. Improvements in this direction would immediately yield improvements in MGS for quasigroups. Furthermore, a constructive membership test would also yield improvements for isomorphism testing of $O(1)$-generated quasigroups. Note that isomorphism testing of $O(1)$-generated groups is known to belong to $\mathsf{L}$ [Tan13].

### v5

@@ 7.2 MGS for Quasigroups
The proof of Lemma 7.13 shows that for any quasigroup $Q$ and any generating set $S\subseteq Q$, every element $g\in Q$ can be realized with a parse tree of depth $O(\log^{2}|Q|)$. In contrast, Wolf [59] establishes the existence of parse trees of depth $O(\log|Q|)$.
For proving Theorem 7.11 we follow essentially the ideas of [25] (though we avoid introducing the notion of Cayley circuits); however, we have to use a more refined approach by not simply guessing an SLP but the complete information given in Lemma 7.13. Fleischer obtained a $\mathsf{quasiAC^{0}}$ bound for Membership for groups by then showing that the Cayley circuits for this problem can be simulated by a depth-$2$ $\mathsf{quasiAC^{0}}$ circuit. We will instead directly analyze the straight-line programs using an $\exists^{\log^{2}n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ (sequential) algorithm.
Proof 7.16 (Proof of Theorem 7.11).
@@
[…]
Proof 7.17 (Proof of Theorem 7.8).
(a) Let $G$ denote the input quasigroup (of order $n$). First, every quasigroup has a generating set of size $\leq\lceil\log n\rceil$ [43]. Therefore, we start by guessing a subset $X\subseteq G$ of size at most $\leq\lceil\log n\rceil$ (resp. the size bound given in the input). For this, we use $O(\log^{2}n)$ existentially quantified non-deterministic bits ($\exists^{\log^{2}n}$). Furthermore, we also guess some additional information, namely, the sequence $z_{1},\dots,z_{k}\in G$ from Lemma 7.13 and, just as in the proof of Theorem 7.11, the witnesses that this is, indeed, a cube-like sequence. As in the proof of Theorem 7.11 this takes again $O(\log^{2}n)$ existentially quantified bits.
In the next step, we verify whether $X$ actually generates $G$. This is done by checking for all $g\in G$ (universally quantifying $O(\log n)$ bits, $\forall^{\log n}$) whether $g\in\langle X\rangle$, which can be done in $\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{quasiAC^{0}}$ just as in Theorem 7.11. Here it is crucial to note that, as we have already guessed the cube-like generating set in the outer existential block, it suffices to guess $O(\log n)$ bits for the exponents. Finally, as in the proof of Theorem 7.11, we verify in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ using the additional information guessed above that $z_{1},\dots,z_{k}\in G$ is indeed, a cube-like sequence. This concludes the proof of the bound (a) for the decision variant.
(b) To find a minimum-sized generating sequence, we can enumerate all possible generating sets $X$ of size at most $\log n$ in $\mathsf{quasiAC^{0}}\cap\mathsf{DSPACE}(\log^{2}n)$. If we want to compute the minimum generating set, we first, have to find a generating set $X$ and then we have to check whether $X$ is actually of smallest possible size. To do this, in a last step, we use the decision variant from above to check that there is no generating set of size at most $\left|\mathinner{X}\right|-1$ for $G$.
@@
While it is possible to directly reduce MGS to Membership in $\mathsf{quasiAC^{0}}$, we obtain slightly better bounds by instead directly leveraging the Reachability Lemma 7.13. It is also possible to $\mathsf{quasiAC^{0}}$-reduce Quasigroup Isomorphism to Membership. This formalizes the intuition that, from the perspective of $\mathsf{quasiAC^{0}}$, membership testing is an essential subroutine for isomorphism testing and MGS.
This might seem surprising, as in the setting of groups, Membership belongs to $\mathsf{L}$, while MGS belongs to $\mathsf{AC}^{1}(\mathsf{L})$ (Thm. 7.1), yet it is a longstanding open problem whether Group Isomorphism is in $\mathsf{P}$.
@@ 7.3 MGS for Magmas
The decision variant of Minimum Generating Set for commutative magmas (unital or not) is $\mathsf{NP}$-complete under many-one $\mathsf{AC^{0}}$ reductions.
This $\mathsf{NP}$-completeness result helps explain the use of Integer Linear Programming in practical heuristic algorithms for the search version of this problem in magmas, e. g., [33].
For the closely related problem they called Log Generators—given the multiplication table of a binary function (=magma) of order $n$, decide whether it has a generating set of size $\leq\lceil\log_{2}n\rceil$—Papadimitriou and Yannakakis proved that Log Generators of Magmas is $\exists^{\log^{2}n}\mathsf{P}$-complete under polynomial-time reductions [45, Thm. 7]. Our proof uses a generating set of size roughly $\sqrt{n}$, where $n$ is the order of the magma; this is analogous to the situation that Log-Clique is in $\exists^{\log^{2}n}\mathsf{P}$, while finding cliques of size $\Theta(\sqrt{n})$ in an $n$-vertex graph is $\mathsf{NP}$-complete.
Proof 7.20.
@@
Aside from multiplication being commutative (e. g., we also have $X_{i}S=\overline{X}_{i}$, etc.), all other products are defined as 0.
The idea is that the presence of $S_{j,k}$ in a word indicates that clauses $j,j+1,\dotsc,k$ have been satisfied. This interpretation aligns with the multiplication above: viz. $C_{j}X=S_{j,j}$ if $X$ satisfies $C_{j}$. If clauses $j$ through $k$ are satisfied ($S_{j,k}$) and clauses $k+1$ through $\ell$ are satisfied ($S_{k+1,\ell}$), then in fact clauses $j$ through $\ell$ are satisfied ($S_{j,k}S_{k+1,\ell}=S_{j,\ell}$). We use “SS” for all of these as a mnemonic for “satisfied” and also because $S=S_{1,m}$ acts as a “swap” on literals.
@@
Identity elements. We may add a new element $e$, relations $ea=ae=a$ for all $a$, and include $e$ as a generator. When this is done, the reduction queries whether the algebra is generated by $n+m+1$ elements or more than that many, since the element $e$ must be contained in every generating set.
As with most $\mathsf{NP}$-complete decision versions of optimization problems, we expect that the exact version—given a magma $M$ and an integer $k$, decide whether the minimum generating set has size exactly kk—is $\mathsf{DP}$-complete, but we leave that as a (minor) open question.
@@ 8 Conclusion
The analogous result is known for groups, by reducing to the connectivity problem on Cayley graphs. The best known bound for quasigroups is $\mathsf{SAC}^{1}$ due to Wagner [58]. Improvements in this direction would immediately yield improvements in MGS for quasigroups. Furthermore, a constructive membership test would also yield improvements for isomorphism testing of $O(1)$-generated quasigroups. Note that isomorphism testing of $O(1)$-generated groups is known to belong to $\mathsf{L}$ [53].